OpenStudy (anonymous):

Which point best represents √75 on this number line? A. point R B. point S C. point T D. point V

OpenStudy (anonymous):

DO NOT GIVE AN ANSWER!!!! I WANNA SEE IF I CAN FIGURE IT OUT FIRST SO THAT I CAN GET SOMEONE TO CHECK IT!!!

OpenStudy (anonymous):

\[\sqrt{75}=5\sqrt{3}=8.66025403784\]

OpenStudy (tkhunny):

\(\sqrt{64} = 8\) \(\sqrt{81} = 9\) Anything less than 8 or greater than 9 is NOT a candidate.

OpenStudy (anonymous):

Sorry, I was helping my mom. I though it was.... A?

OpenStudy (tkhunny):

Point "R" is the only one that is even close.

OpenStudy (anonymous):

That's what I thought. I just wanted to make sure. :) Thanks! :):):)